OpenStudy (anonymous):

can i get some help with exponent rules? My teacher wants me to memorize them.

OpenStudy (anonymous):

x1 = x x0 = 1 x-1 = 1/x xmxn = xm+n xm/xn = xm-n (xm)n = xmn (xy)n = xnyn (x/y)n = xn/yn

OpenStudy (anonymous):

x1 = x^1 sorry did it too quick

OpenStudy (anonymous):

no problem:)

OpenStudy (anonymous):

oh and can you please explain the rules instead of just typing them out. their easier to memorize that way, thanks.

OpenStudy (anonymous):

I mean the main ones you need to remember is that x^m * x^n = x^(m+n) and x^m / x^n = x^(m-n)

OpenStudy (anonymous):

What do you mean explain?>

OpenStudy (anonymous):

i mean instead of giving me the formula, explain what it means like instead of 0^x=1 say 0 to the x equals one. explain the variables basically.

OpenStudy (anonymous):

They are just rules.. Like x^2 can be written as x*x and x^3 can be written as x*x*x, so if you times x^2 and x^3, you get x*x*x*x*x, which is equal to x^5...

OpenStudy (anonymous):

i get that,

OpenStudy (anonymous):

x^0 = 1 is just a definition.. there isn't really a proof out there that proves it.

OpenStudy (anonymous):

thanks.

OpenStudy (anonymous):

i get that,

OpenStudy (anonymous):

k